OpenStudy (anonymous):

Jan is 5 years older than Terry. In 3 years, Terry will be 3/4 as old as Jan will be. What is Terry's age now?

OpenStudy (anonymous):

thats a very inappropriate question to ask someone

OpenStudy (anonymous):

Let J be Jan's age, T be Terries age. What is the equation represented by the first statement "Jan is 5 years older than Terry"

OpenStudy (anonymous):

J=T+5

OpenStudy (anonymous):

Correct. Now, what is the equation represented by the statment "In 3 years Terry will be 3/4 as old as Jan"

OpenStudy (anonymous):

That's the part I am having trouble with.??? T+3=(3/4)J

OpenStudy (anonymous):

Very nearly, but recall that in 3 years both will be 3 years older than they are now.

OpenStudy (anonymous):

T+3=(3/4)J+3 ????

OpenStudy (anonymous):

Terry's age in 3 years = T+3 Jan's age in 3 years = J+3 Terry's age in 3 years is 3/4 Jan's age in 3 years \(\implies T+3 = \frac{3}{4}(J+3)\)

OpenStudy (anonymous):

Okay, I get it. I can finish it from here. Thanks so much!!!

OpenStudy (anonymous):

Of course!